Books cost 40 cents and pamphlets 5 cents at the book sale. If Mr. Jones spent $4.15 at the book sale and purchased 20 more pamphlets than he did books, how many pamphlets did he buy?
Answer by jorel1380(3719)

p=b+20
40b+5(b+20)=415
40b+5b+100=415
45b=315
b=315/45
b=7
p=27
Mr. Jones bought 7 books and 27 pamphlets..